# Oblack Owhite m #3

### v1

@@ C.1 More demonstrations of harmful records
We examine the bottom records w.r.t $f^{\text{return}}$ in different training rounds $k$ and present the results in Fig. 8. (Results in the main paper, Fig. 3(a), corresponds to $k=5$ here.)
@@
C.2 Quantifying phase change via weighted graph roughness analysis
@@
We study another environment Acrobot, investigating the phase change and measuring the roughness metric across rounds. The results are presented in Fig. 9. We observe a consistent trend of the three phases, aligned with the findings discussed in Sec. 4.2.
@@
C.3 Additional results for single-round intervention
  Influence-guided intervention   Random drop
@@ C.3 Additional results for single-round intervention
Fig. 10 (as an extension of Fig. 5) presents the results of single-round interventions in four environments, additionally comparing with the random baseline that discards a similar amount of records.
We discuss several key takeaways: (1) Influence-guided intervention mostly leads to performance gains, while random drop mostly leads to performance degradation. (2) When standard PPO fails to improve (e.g. a dip at round $k=9$ in Highway; see Fig. 6), the attribution signal can become unreliable, producing negative $\Delta$ return (see Fig. 10 at $k=9$ in Highway), leading occasionally to interventions that fail to bring any improvement. However, as long as PPO’s overall trend is upward, our intervention can effectively purify the learning and and drive net improvement over the full run.
C.4 Advantage-based heuristic
@@
Sec. 4.1 characterizes the properties of the bottom harmful records—sign mismatch and large magnitude errors. Inspired by these findings, we design the following two heuristics for experience filtering:
  • Heuristic 1: We discard records with opposite signs for $\bar{A}$ and $\hat{A}$. Among these records, we sort them by $|\bar{A}-\hat{A}|$ and discard the top $p\%$ records with the largest error.
@@ Results.
Fig. 11 compares the two advantage-based heuristics against IIF and standard training in FrozenLake and MiniGrid.
@@
Figure 11: Test returns over training rounds for the two advantage-based heuristics, compared with IIF and standard PPO. Results are averaged over three random seeds.
@@
C.5 TD error based heuristic
@@
Prioritized Experience Replay (PER) [Schaul et al., 2016] demonstrate that reweighting transitions in proportion to their temporal-difference (TD) error accelerates learning and improves performance in off-policy methods. TD error serves as a useful heuristic, indicating how “surprising” or “important” a transition is for updating the value function. While PPO is an on-policy method that typically uses a smaller, on-policy rollout buffer rather than a large replay buffer like those in off-policy algorithms, the core idea of focusing learning on more impactful experiences remains relevant. Inspired by PER, we investigate integrating a TD error based reweighting mechanism into the PPO training process to prioritize samples within its rollout buffer.
@@ Results.
We evaluate the performance of the TD error based reweighting heuristic by comparing it against our IIF and standard PPO on FrozenLake and LunarLander. Fig. 12 presents the test returns over training rounds for these approaches.
@@
In FrozenLake, a simple environment, both TD error and IIF accelerate convergence, reaching optimal return sooner. The TD error heuristic nearly matches IIF’s speed, confirming that large TD errors align well with truly useful transitions when the state-action space is small and reward structure simple.
In contrast, in the more complex LunarLander, the TD error heuristic degrades performance: it learns more slowly than even standard PPO and exhibits greater variance. Although this heuristic succeeds in PER, we comment that there are intrinsic differences in the off-policy scenario where PER was proposed and evaluated, vs. the on-policy scenario (e.g., PPO) we study in this paper (Fig. 1). PER applies the TD error heuristic on a vast, diverse buffer. However, in PPO, raw TD errors mix estimator noise with true signal; PPO’s small, fresh, on-policy batches exacerbate that noise; Our influence scores, in comparison, appears more robust in such scenarios.
C.6 IIF performance under various filtering percentages
@@
Fig. 13 shows the test returns over training rounds for IIF with varying $p$’s compared to baselines. We additionally quantify their efficiency using two metrics: $SE_{\text{ave}}$ and $SE_{\text{peak}}$ (introduced in Sec. 5.2). We summarize these efficiency statistics in Table 2.
@@
  • Discarding all negative records ($p=100\%$) is suboptimal. As shown in Figure 13, setting $p=100\%$ leads to suboptimal final performance, slower learning progress (also reflected in Table 2), and instability in training. This observation aligns with the concept of non-additivity of sample influence [Hu et al., 2024].
@@
Based on these findings, for our main experiments (see Sec. 5.2) we choose the specific filtering percentages to reflect the optimal configuration per environment. We use $p=50\%$ for FrozenLake, Acrobot, Highway; $p=12.5\%$ for MiniGrid, LunarLander; and $p=6.25\%$ for BipedalWalker.
C.7 Runtime for experiments on traditional RL environments
We report the runtime for experiments on traditional RL environments in Table 3.
@@ C.7 Runtime for experiments on traditional RL environments
Finally, we report $RT_{\text{peak}}$ (also presented in Fig. 6(b)), calculated as the reduced percentage of wall clock time for IIF to match standard PPO. In summary, IIF presents a 29%-67% reduction in runtime, effectively speeding up learning.
@@
Per-round runtime Influence calc 0.11 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.02
@@
Per-round runtime Influence calc 0.13 $\pm$ 0.02 0.13 $\pm$ 0.01 0.12 $\pm$ 0.01
@@
C.8 Comparing two target functions for RLHF
In the main text (Sec. 5.3), we introduced two target functions for RLHF: the standard one $f^{\text{return}}$, and an adapted sequence-level objective $f^{\text{seq}}$. Here we show the comparison of the two in Fig. 14.
(a) Training reward ($\uparrow$) (b) Test toxicity ($\downarrow$)
@@ C.8 Comparing two target functions for RLHF
C.9 A breakdown of runtime for the RLHF experiments
Table 4 breaks down the wall-clock time (in seconds) for each component of one RLHF training round, under standard PPO and our IIF. The overhead of influence calculation in IIF is significantly offset by reduced optimization time, leading to a 2×2$\times$2 × speedup per round.
Beyond this per-round saving, IIF requires fewer rounds to achieve comparable performance with standard PPO (requiring 32.75% $\pm$ 1.52% of training rounds, taking up 16.82% $\pm$ 1.32% of runtime combined with per-round speedup). Furthermore, IIF reaches convergence to a higher reward faster as well (requiring 48.51% $\pm$ 2.44% of training rounds, taking up 24.90% $\pm$ 0.80% of wall-clock time). This marks a 4×4$\times$4 × overall speedup plus performance improvement compared to standard PPO.
@@ C.9 A breakdown of runtime for the RLHF experiments
Appendix D Compute resources
@@
All runtime results reported in Sec. C.7 were measured on Machine 1; all runtime results in Sec. C.9 were measured on Machine 2.

### v2

@@ B.1 More demonstrations of harmful records
We examine the bottom records w.r.t $f^{\text{return}}$ in different training rounds $k$ and present the results in Fig.˜8. (Results in the main paper, Fig.˜3(a), corresponds to $k=5$ here.)
@@
Harmful records in complex environments.
We look into two complex environments. In BipedalWalker (locomotion), our analysis reveals bottom records where the agent was incorrectly penalized with a large negative advantage for executing a successful recovery move (e.g., applying corrective torque with a deeply bent knee ($\sim$35°) during landing or push‑off). (We omit the visualizations for this environment as it does not conveniently support rendering given status vectors; the above analysis is done based on direct analysis of values in status vectors.) In Pong (Atari), we find that bottom records filtered by IIF consist of uninformative transitions (the ball being out of play or already moving away from the agent) that receive (inaccurately) high advantage estimates. By filtering out these samples, IIF achieves significant improvement in training efficiency. These results show that 1) bottom records feature inaccurate advantage estimates; 2) IIF is effective, holding generally across different environments. Examples are shown in Fig.˜9.
Figure 9: Bottom records for the Pong. The top and middle figures correspond to the case where the ball it out of play. The bottom figure corresponds to the case where the ball is moving away from the agent. (Note that in Pong, the ego agent is the one on the right.)
B.2 Quantifying phase change via weighted graph roughness analysis
@@
We study another environment Acrobot, investigating the phase change and measuring the roughness metric across rounds. The results are presented in Fig.˜10. We observe a consistent trend of the three phases, aligned with the findings discussed in Sec.˜4.2.
@@
B.3 Additional results for single-round intervention
  Influence-guided intervention
  Random drop
@@ B.3 Additional results for single-round intervention
Fig.˜11 (as an extension of Fig.˜5) presents the results of single-round interventions in four environments, additionally comparing with the random baseline that discards a similar amount of records.
We discuss several key takeaways: (1) Influence-guided intervention mostly leads to performance gains, while random drop mostly leads to performance degradation. (2) When standard PPO fails to improve (e.g. a dip at round $k=9$ in Highway; see Fig.˜6), the attribution signal can become unreliable, producing negative $\Delta$ return (see Fig.˜11 at $k=9$ in Highway), leading occasionally to interventions that fail to bring any improvement. However, as long as PPO’s overall trend is upward, our intervention can effectively purify the learning and and drive net improvement over the full run.
B.4 Advantage-based heuristic
@@
Sec.˜4.1 characterizes the properties of the bottom harmful records—sign mismatch and large magnitude errors. Inspired by these findings, we design the following two heuristics for experience filtering:
[…]
@@ Results.
Fig.˜12 compares the two advantage‐based heuristics against IIF and standard training in FrozenLake and MiniGrid.
@@
Figure 12: Test returns over training rounds for the two advantage‐based heuristics, compared with IIF and standard PPO. Results are averaged over three random seeds.
@@
B.5 TD error based heuristic
@@
Prioritized Experience Replay (PER) [Schaul et al., 2016] demonstrate that reweighting transitions in proportion to their temporal‐difference (TD) error accelerates learning and improves performance in off‐policy methods. TD error serves as a useful heuristic, indicating how “surprising” or “important” a transition is for updating the value function. While PPO is an on-policy method that typically uses a smaller, on-policy rollout buffer rather than a large replay buffer like those in off-policy algorithms, the core idea of focusing learning on more impactful experiences remains relevant. Inspired by PER, we investigate integrating a TD error based reweighting mechanism into the PPO training process to prioritize samples within its rollout buffer.
@@ Results.
We evaluate the performance of the TD error based reweighting heuristic by comparing it against our IIF and standard PPO on FrozenLake and LunarLander. Fig.˜13 presents the test returns over training rounds for these approaches.
@@
In FrozenLake, a simple environment, both TD error and IIF accelerate convergence, reaching optimal return sooner. The TD error heuristic nearly matches IIF’s speed, confirming that large TD errors align well with truly useful transitions when the state‐action space is small and reward structure simple.
In contrast, in the more complex LunarLander, the TD error heuristic degrades performance: it learns more slowly than even standard PPO and exhibits greater variance. Although this heuristic succeeds in PER, we comment that there are intrinsic differences in the off-policy scenario where PER was proposed and evaluated, vs. the on-policy scenario (e.g., PPO) we study in this paper (Fig.˜1). PER applies the TD error heuristic on a vast, diverse buffer. However, in PPO, raw TD errors mix estimator noise with true signal; PPO’s small, fresh, on-policy batches exacerbate that noise; Our influence scores, in comparison, appears more robust in such scenarios.
B.6 IIF performance under various filtering percentages
@@
Fig.˜14 shows the test returns over training rounds for IIF with varying $p$’s compared to baselines. We additionally quantify their efficiency using two metrics: $SE_{\text{ave}}$ and $SE_{\text{peak}}$ (introduced in Sec.˜5.2). We summarize these efficiency statistics in Table˜2.
@@
  • Discarding all negative records ($p=100\%$) is suboptimal. As shown in Figure 14, setting $p=100\%$ leads to suboptimal final performance, slower learning progress (also reflected in Table˜2), and instability in training. This observation aligns with the concept of non-additivity of sample influence [Hu et al., 2024].
@@
Based on these findings, for our main experiments (see Sec.˜5.2) we choose the specific filtering percentages to reflect the optimal configuration per environment. We use $p=50\%$ for FrozenLake, Acrobot, Highway; $p=12.5\%$ for MiniGrid, LunarLander; and $p=6.25\%$ for BipedalWalker.
B.7 Evaluating IIF with the Adam optimizer
Our main experiments in traditional RL environments are conducted using the SGD optimizer (see Sec.˜A.2). Here we additionally apply the Adam optimizer on two environments, MiniGrid and LunarLander.
We report the test return in Fig.˜15, and sample efficiency and runtime metrics in Table˜3. One observation is that IIF gains less with Adam compared to SGD in MiniGrid, whereas the trend is reversed for LunarLander (see Fig.˜6 for reference). This is partly because Adam significantly speeds up training compared to SGD in MiniGrid (and thus reduces the room of improvement), but less so in LunarLander.
Figure 15: Test returns over rounds for IIF vs. the standard training baseline, when using the Adam optimizer. Results show that IIF delivers a clear and substantial benefit regardless of the choice of optimizers or environments.
Table 3: Sample efficiency and runtime comparisons when using the Adam optimizer.
MiniGrid LunarLander
$SE_{\text{ave}}\ (\uparrow)$
24.1% $\pm$ 1.4% 46.7% $\pm$ 4.5%
$SE_{\text{peak}}\ (\uparrow)$
13.3% $\pm$ 3.1% 62.2% $\pm$ 5.0%
$RT_{\text{peak}}\ (\uparrow)$ 18.5% $\pm$ 1.0% 65.9% $\pm$ 3.2%
B.8 Statistical significance of final performance gains
We compute the 95% confidence interval (CI) for the performance gain of IIF over the standard baseline (as shown in Fig.˜6(a)). Concretely, we compute half-width = $t_{0.957,4}\times SE=2.776\times SE$. Results in Table˜4 confirm a statistically significant improvement in the performance gain.
Table 4: 95% confidence interval (CI) for the performance gain of IIF over the standard baseline across 5 random seeds.
MiniGrid LunarLander BipedalWalker
95% CI [0.04, 0.33] [22.54, 130.52] [24.40, 75.99]
B.9 Runtime for experiments on traditional RL environments
We report the runtime for experiments on traditional RL environments in Table˜5.
@@ B.9 Runtime for experiments on traditional RL environments
Finally, we report $RT_{\text{peak}}$ (also presented in Fig.˜6(b)), calculated as the reduced percentage of wall clock time for IIF to match standard PPO. In summary, IIF presents a 29%-67% reduction in runtime, effectively speeding up learning.
@@
Per-round
runtime
Influence calc 0.11 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.02
@@
Per-round
runtime
Influence calc 0.13 $\pm$ 0.02 0.13 $\pm$ 0.01 0.12 $\pm$ 0.01
@@
B.10 Difficulty based heuristic
Inspired by the difficulty-based filtering (e.g., pass@k) primarily used to improve LLM Reasoning (RLVR) in GRPO [Yu et al., 2025, Bae et al., 2025], we develop a difficulty-based filtering approach for PPO. Concretely, we use reward as a proxy for difficulty and filter records receiving top and bottom rewards. However, this heuristic performs worse than random because it systematically removes data with both highest and lowest influence scores, thereby harming the learning process. This finding aligns with our results in Sec.˜B.5 for traditional RL, where an analogous heuristic using TD error as a proxy for difficulty also proved ineffective. Therefore, our evidence shows that while valid for GRPO, difficulty-based filtering is an ineffective heuristic for PPO.
B.11 Comparing two target functions for RLHF
In the main text (Sec.˜5.3), we introduced two target functions for RLHF: the standard one $f^{\text{return}}$, and an adapted sequence-level objective $f^{\text{seq}}$. Here we show the comparison of the two in Fig.˜16.
(a) Training reward ($\uparrow$)
(b) Test toxicity ($\downarrow$)
@@ B.11 Comparing two target functions for RLHF
B.12 A breakdown of runtime for the RLHF experiments
Table 6 breaks down the wall-clock time (in seconds) for each component of one RLHF training round, under standard PPO and our IIF. The overhead of influence calculation in IIF is significantly offset by reduced optimization time, leading to a $2\times$ speedup per round.
Beyond this per-round saving, IIF requires fewer rounds to achieve comparable performance with standard PPO (requiring 32.75% $\pm$ 1.52% of training rounds, taking up 16.82% $\pm$ 1.32% of runtime combined with per-round speedup). Furthermore, IIF reaches convergence to a higher reward faster as well (requiring 48.51% $\pm$ 2.44% of training rounds, taking up 24.90% $\pm$ 0.80% of wall-clock time). This marks a $4\times$ overall speedup plus performance improvement compared to standard PPO.
@@ B.12 A breakdown of runtime for the RLHF experiments
Appendix C Compute resources
@@
All runtime results reported in Sec.˜B.9 were measured on Machine 1; all runtime results in Sec.˜B.12 were measured on Machine 2.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We have tried our best to ensure that the abstract and introduction accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations are discussed in Sec.˜7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide detailed information on the experimental setups in Appendix˜A. Our code is also also publicly available at https://github.com/LDAORL/LDA-ORL.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our code is publicly available at https://github.com/LDAORL/LDA-ORL.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The details of the experiments are discussed in Appendix˜A.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We use 3 random seeds for all experiments; we include error bars in all reported results (Figs.˜6, 5 and 7) in the main paper as well as more results in Appendix˜B (Figs.˜11, 12, 13, 14, 2, 5 and 6).
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The information on the compute resources is provided in Appendix˜C.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS Code of Ethics and confirm that the research conducted in this paper adheres to its principles.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The paper conducts fundamental research aimed at understanding the role of data in online RL, and leverages this understanding to improve RL training. We do not anticipate any immediate societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not pose such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited the RL environments / datasets, models, code frameworks, and included their licenses in Appendix˜A.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The paper does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.